OpenStudy (anonymous):

Find the complex zeros of the following: x^3 + 216 Write in factored form.

OpenStudy (anonymous):

Does this equation equal anything?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Refresh my memory. A "complex zero" is of the form a +bi, which makes the given equation =0, correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

216 is 6 cubed. So the zeros are going to be six times the zeros of x^3 + 1=0. Can you do that problem?

OpenStudy (anonymous):

Since you want the zeros, we should set the equation equal to 0 and solve by factoring. we get: \[x^{3}+216 = (x+6)(x^{2}-6x+36)=0\] Now we set both of those terms equal to 0, but the left term isnt going to produce a complex root: \[x^{2}-6x+36 = 0 \Rightarrow x = \frac{6\pm \sqrt{36-4(36)}}{2} = 3 \pm 3i \sqrt{3}\] If you want me to show more work let me know :)

myininaya (myininaya):

(x+6)(x^2-6x+36)=0